if 3 is subtracted from a number and the remainder is multiplied by 13,the product becomes 91.find the number

Answer by Theo(13342)

x - 3 is the remainder when you subtract 3 from x.
13 * (x-3) = 91
simplify to get 13x - 39 = 91
add 39 to both sides of the equation to get 13x = 130
divide both sides of the eqaution by 13 to get x = 10

when x = 10, x-3 = 7
7 * 13 = 91